How long will it take to make your investment double if the expected annual return is 15%?

5 years
6 years
4 years
3 years
Answer #1

Present Value = Future Value / (1+r)^n

Present Value be X

Future Value be 2X

X = 2X / (1+0.15)^n

1 = 2 / (1+0.15)^n

1.015^n = 2

Taking log both sides

n * Log1.15 = Log 2

n * 0.06069784035 = 0.30102999566

n = 0.30102999566 / 0.0646604224

n = 4.65 OR 5 years (approx)

Option A is correct.